# StoneDetector: Conventional and versatile code clone detection for Java

### v1

@@ Abstract
Copy & paste is a widespread practice when developing software and, thus, duplicated and subsequently modified code occurs frequently in software projects. Since such code clones, i.e., identical or similar fragments of code, can bloat software projects and cause issues like bug or vulnerability propagation, their identification is of importance. In this paper, we present the StoneDetector platform and its underlying method for finding code clones in Java source and Bytecode. StoneDetector implements a conventional clone detection approach based upon the textual comparison of paths derived from the code’s representation by dominator trees. In this way, the tool does not only find exact and syntactically similar near-miss code clones, but also code clones that are harder to detect due to their larger variety in the syntax. We demonstrate StoneDetector’s versatility as a conventional clone detection platform and analyze its various available configuration parameters, including the usage of different string metrics, hashing algorithms, etc. In our exhaustive evaluation with other conventional clone detectors on several state-of-the-art benchmarks, we can show StoneDetector’s performance and scalability in finding code clones in both, Java source and Bytecode.
@@ 1 Introduction
Code clones [1] are identical or similar fragments of code that, e.g., arise due to copy-and-paste practices in software development. Since such practices are quite common, code clones can be frequently found, as empirical studies on code duplication have shown in the past [2, 3]. While code cloning may initially speed up software development or allow for reusing working code, it also introduces long-term challenges: Duplicated code inflates code bases and increases the maintenance burden, as developers have to manually track and update all copies for fixing bugs or implementing code changes. The effects of code cloning are thus controversial [4, 5] and detecting code clones is a vital step toward maintaining code quality, reducing technical debt, and improving program comprehension.
@@ 10 FileInputStream fis = new FileInputStream(f);
The presented clone detection method is implemented in the StoneDetector platform, which is publicly available [^1] https://stonedetector.fmi.uni-jena.de/ and allows for finding code clones in Java source and Bytecode. Using the platform enabled us to perform an in-depth evaluation of the method’s performance on various state-of-the-art benchmarks, including BigCloneBench[12, 13], Google Code Jam, Project CodeNet [15], and GPTCloneBench [16]. As a result, we found that our clone detection method and its implementation in the StoneDetector platform allow for finding exact and near-miss code clones in Java source code at competitive performance with other conventional code clone detectors, including NiCad [14], iClones [17], SourcererCC [18], Deckard [19], CCAligner [10], NIL [20], and Oreo [11]. Additionally, StoneDetector in particular excels at finding source code clones with larger syntactical variance at high precision. StoneDetector also scales to large codebases comprising multiple 100 million lines of code and features runtimes comparable to the state of the art. Furthermore, we have experimented with different configuration parameters for clone detection, e.g., threshold value, code clone size, applied string metrics, and hashing algorithms, etc., and thereby show the versatility of the StoneDetector platform besides identifying its optimal configuration. Eventually, we analyzed the method’s potential in also finding code clones in compilation artifacts like Java Bytecode, in cases where the source code is unavailable in the first place. In summary, the contributions of this paper are:
@@
    We demonstrate the feasibility of finding code clones in Java code fragments by encoding and textually comparing paths of the code fragments’ dominator trees.
@@
    We provide the publicly available StoneDetector platform with an implementation of our code clone detection method as a versatile conventional code clone detector and evaluate its performance and scalability.
@@
This work is based upon the authors’ prior conference paper on code clone detection presented at the 2021 IEEE International Conference on Software Maintenance and Evolution [21] and extends it in various ways. First, an analysis of the usage of different string metrics within StoneDetector has been added to the paper, including distance metrics like Levenshtein, Needleman-Wunsch, Hamming distance, and longest common subsequence (LCS)-based metric. Among others, the analysis revealed the versatility of StoneDetector as a clone detection framework: For example, using LCS provided the best tradeoff in terms of recall, precision, and scalability. In contrast, using the Hamming distance yielded a very fast clone detector with impaired performance in finding harder-to-detect Type 3/4 code clones. Second, we now derive the optimal configuration of StoneDetector for clone detection by means of an in-depth empirical analysis of its several configuration parameters (threshold value for path comparison, etc.). Furthermore, expanding the approach for detecting large-gap and subclones, as presented in our previous conference paper [21], we added locality-sensitive hashing (LSH) and a modified LCS metric to StoneDetector, which resulted in performance gains for these types of code clones. We now also include a more exhaustive analysis of StoneDetector’s performance in finding harder-to-detect Type 3/4 code clones by including three more state-of-the-art benchmarks for code clone detection besides BigCloneBench, i.e., Google Code Jam, Project CodeNet, and GPTCloneBench, and two more clone detectors, i.e., iClones and NIL, in the evaluation. Moreover, we also extended our previous approach in Amme et al. [21] for clone detection on stack-based and register-based Java Bytecode representations, such that the StoneDetector platform can be used for both finding code clones in Java source code as well as Java Bytecode.
@@
The rest of the paper is structured as follows: Our clone detection approach and its underlying code clone detection method based on encodings of paths in the dominator tree and string metrics for path comparison are described in detail in Sect. 2. Sect. 3 introduces the StoneDetector platform’s architecture and implementation. In Sect. 4, we present the in-depth evaluation of our approach on various benchmarks for Java code clone detection, including the comparison with state-of-the-art clone detectors and the thorough analysis of the StoneDetector platform’s various configuration parameters. Related work on code clone detection is discussed in Sect. 5 and Sect. 6 finally concludes the paper.
2 Clone Detection using Dominator Trees
2.1 Control flow and Dominator Trees
A control flow graph $G$ is a directed graph defined as $G=(N,E,s)$ with $(N,E)$ denoting the graph and $s\in N$ the start node [22]. We encode the instructions of a code fragment using the graph’s nodes. Note that we here use a single node for each instruction instead of using basic blocks. An edge $(k,l)\in E$ in a control flow graph connects two nodes $k$ and $l$ iff the instruction denoted by $l$ can be immediately executed after the instruction denoted by $k$. Each node $n\in N$ is then reachable via a path from the start node $s$, i.e., a sequence of consecutive edges $e\in E$ that begin at $s$ and end at $n$. An example of a control flow graph representing the lower code fragment shown in Fig. 2, i.e., function zipDir, can be found in Fig. LABEL:fig2 (a).
A dominator tree [22] is a tree and models a certain control flow property of code fragments, specifically, which instruction precedes the other under all possible executions of the respective code fragment. Thus, let the triple $G=(N,E,s)$ denote a certain control flow graph. A node $l\in N$ dominates a node $k\in N$ iff $l$ occurs on all paths from the start node $s$ to node $k$. Accordingly, every node $n\in N$ dominates itself by definition. A node $l\in N$ furthermore strictly dominates another node $k\in N$ iff $l$ dominates $k$ and $l\neq k$. Node $l$ is also immediately dominating $k$ iff $l$ strictly dominates $k$ and there exists no node $m\in N$, such that $m$ strictly dominates $k$ and $l$ strictly dominates $m$. The dominator tree for the control flow graph $G$ is then a directed graph $D=(N,E^{*})$ with $E^{*}=\{(l,k)\,|\,l\mbox{ immediately dominates }k\}$. An example of a dominator tree is visualized in Fig. LABEL:fig2 (b). As can be seen, the shown dominator tree corresponds to the control flow graph in Fig. LABEL:fig2(a). Note that in the dominator tree, each node is reachable from the start node, i.e., the tree’s root, via a single unambiguous path. This is in contrast to the given control flow graph. Also note that there are no backward edges encoding loop edges, and therefore cycles, in the dominator tree, since the loop body cannot dominate the loop condition according to the above definition.
We use an abstract, text-oriented encoding of instructions for our clone detection method, yielding the abstract dominator tree which is depicted in Fig. LABEL:fig2 (c). Therein, instructions and their respective operands are denoted in preorder notation. Furthermore, the distinct variable and field names, as well as literals, are not distinguished and mapped to a single joint symbol, respectively. Thus, all variable names are subsumed by symbol $V$, literals by symbol $L$, and so on, much like in the clone detector NiCad [9]. For example, consider instruction i = 0 which is encoded by [=,V,L] in the abstract dominator tree (cf. Fig. LABEL:fig2). Function calls are optionally represented using a constant pool. A constant pool access is denoted by the character # and the respective function’s entry in the constant pool, as, e.g., in [CALL,#6,V] for instruction zos.putNextEntry(anEntry) (again, cf. Fig. LABEL:fig2). By using abstract dominator trees, we are able to more easily match code fragments that do not completely align with each other syntactically, in particular in the case of Type 2 and part of Type 3/4 code clone deviations.
2.2 Finding Similar Code Fragments
Our approach is similar to the method proposed by Alon et al. [24], where they use a linearization of abstract syntax trees based on paths that start and end at distinct leaf nodes of an abstract syntax tree. Though, in contrast, in our method, given a dominator tree $D=(N,E^{*})$, we consider all paths starting at the tree’s root node $s\in N$ and ending at a respective leaf node $l\in\{n\in N\,|\,\not\exists(n,m)\in E^{*}\}$. The resulting set of these paths can then be used to unambiguously represent the dominator tree $D$. A single path is encoded by simply concatenating the (abstract) instructions denoted by the path’s nodes into a sequence of instructions. For example, considering the abstract dominator tree in Fig. LABEL:fig2 (c), the path from the start node to the rightmost leaf node is encoded by the sequence:
[…]
Let $D$ be a dominator tree, $\mathit{paths}_{D}$ the set of its paths, and $\mathit{leaves}_{D}$ the set of its leaf nodes in the following. The set of paths unambiguously representing $D$ is given by: $D_{\mathit{set}}=\{s\rightarrow\cdots\rightarrow n_{k}\,|\,s\cdots n_{k}\in\mathit{paths}_{D},n_{k}\in\mathit{leaves}_{D}\}$. We call the set $D_{\mathit{set}}$ the description set of dominator tree $D$. In the case of the abstract dominator tree depicted in Fig. LABEL:fig2 (c), we have four paths starting at the start node and ending in the leftmost, the left inner, the right inner, and the rightmost leaf node, respectively. Accordingly, the description set shown in Fig. LABEL:fig2 (d) contains the four instruction sequences representing the paths.
2.3 Metrics for Path Comparison
The Hamming distance [25] is a basic string metric, which is defined on text strings with fixed length. This distance measures the minimum number of substitutions to change one string into another and thus yields the number of deviating characters in two given strings. Therefore, as a limitation, the text strings are only compared up to the length of the respective shorter string, and the number of remaining characters in the longer string is merely added to the so far calculated distance measure. Note that the conventional Hamming distance therefore tends to penalize paths with different lengths.
Example:
  • Code fragment 1: for (int j=0; j<<m; j++) { acc=acc-j; proc(acc); }
  • $p=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}\rightarrow\texttt{[POSTINC,V]}$
  • Code fragment 2: for (int l=0; l<<m; l++) { acc=acc+l; }
  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\rightarrow$ [POSTINC,V]
Hamming:
  • 1.
    Substitute [=,V,+,V,V] with [=,V,-,V,V]
    $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[POSTINC,V]}$
  • 2.
    Substitute [POSTINC,V] with [CALL,#4,V]
    $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}$
  • Hamming distance $\delta_{\mathit{H}}(p,q)=3=2+1$ because $p$ has one more character than $q$.
Levenshtein:
  • 1.
    Substitute [=,V,+,V,V] with [=,V,-,V,V]
    $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[POSTINC,V]}$
  • 2.
    Insert [CALL,#4,V]
  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[\mbox{CALL},\#4,V]}\rightarrow\texttt{[POSTINC,V]}$
  • Levenshtein distance $\delta_{\mathit{L}}(p,q)=2$.
Needleman-Wunsch:
  • $p=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\_\_\_\_\_\_\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}$ $\rightarrow\texttt{[POSTINC,V]}$
  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\rightarrow\_\_\_\_\_\_\rightarrow\_\_\_\_\_\_\rightarrow\texttt{[POSTINC,V]}$
  • Needleman-Wunsch-based distance $\delta_{\mathit{NW}}(p,q)=max(|p|-score(p,q),|q|-score(p,q))=2$, where $score(p,q)=3$.
Longest Common Subsequence (LCS):
  • $p=\texttt{[=,V,L]}\rightarrow\texttt{[COND,\textless,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}\rightarrow\texttt{[POSTINC,V]}$
  • $q=\texttt{[=,V,L]}\rightarrow\texttt{[COND,\textless,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\ \rightarrow\texttt{[POSTINC,V]}$
  • $z=\texttt{[=,V,L]}\rightarrow\texttt{[COND,\textless,V,V]}\rightarrow\texttt{[POSTINC,V]}$
  • LCS-based distance $\delta_{\mathit{LCS}}(f,g)=max(|p|-|z|,|q|-|z|)=2$.
Figure 4: Distance metrics for calculating path similarity
Another metric that has its origins in the alignment of DNA sequences is the Needleman-Wunsch algorithm [27]. The algorithm computes an alignment of two given text strings that is optimal with respect to a certain scoring scheme using two steps. In the first step, a similarity matrix of the two text strings is generated based upon the scoring scheme, which defines weights for character matches, mismatches, and gaps. In the second step, the optimal string alignment with maximal total score is computed by performing a backward traversal on this similarity matrix.
The application of string metrics to the comparison of paths in abstract dominator trees is illustrated in Fig. 4 by way of an example. In the upper part, two code fragments are shown, conjoined with two paths $p$ and $q$ from their abstract dominator trees. As can be seen below, the (abstract) instructions take on the role of characters in the visualized calculations of string metrics. In the alignment provided for the Needleman-Wunsch algorithm, vertical lines between nodes represent matches and horizontal lines represent inserted gaps. For the value of $store(p,q)$, a scheme is used in which the weights for matches are set to $1$, for mismatches to 1-1, and for gaps to 0.
The Longest Common Subsequence (LCS) algorithm [28, 9] is a well-known tool for comparing the contents of files, used for example in the Unix diff utility or in the revision control system git. Besides, clone detectors like NiCad are relying on the LCS algorithm as well. In our case, the algorithm is applied to paths instead of lines of code, such that the longest common subpath is computed, e.g., $\texttt{[=,V,L]}\rightarrow\texttt{[COND,<,V,V]}\rightarrow\texttt{[POSTINC,V]}$ for paths $p$ and $q$ in Fig. 4. This longest common subpath then forms the basis for defining the metric $\delta$ for two paths of dominator trees, which equals the maximum number of unique nodes for both paths.
Using the original LCS-based metric, though, can become problematic in cases of large-gap code clones [10], e.g., code clones where one code fragment is embedded into the other. The LCS metric then has issues in recognizing the similarity of the respective paths of the code fragments’ dominator trees, since the paths differ too much in their length. For this reason, a modified variant of the LCS-based metric can be used, which performs additional similarity checking in situations where two paths are not considered similar by the original LCS-based metric and analyzes whether or not the derived LCS distance value is close to the length of a respective smaller path. On a technical note, code fragments consisting of only one basic block and thus path are then considered as code clones if the calculated LCS value corresponds to a value lower than 30% of the number of nodes of the smaller code fragment. Furthermore, a path $p$ is also considered a subpath of a path $q$ if the calculated LCS value of $p$ and $q$ is less than 10% of the length of $p$. Finally, the smaller of both paths is required to have a minimum length of 10 nodes.
Note that there exist other string metrics as well, besides the ones discussed above, which have also been implemented in the StoneDetector platform. Among these are, e.g., N-gram distance, cosine distance, Jaccard similarity. For the sake of brevity, we here limit ourselves to the subset of metrics that we use in the evaluation in Sect. 4. For a more comprehensive discussion of string metrics, we refer, e.g., to the survey [29].
2.4 Fingerprints and Hashing
However, using string metrics on the plain paths of (abstract) dominator trees for path comparison is rather inefficient due to the need for node matching, i.e., matching equal instructions, multiple times for a single comparison of two paths. Therefore, to minimize the costs, we conduct the node matching on fingerprints instead, which encode the respective nodes’ instructions. Such a fingerprint is simply the hashed value of an instruction, for which the string metrics are adjusted accordingly. There are apparently various hashing algorithms available for this purpose. For example, the MD5 algorithm [30] provides a widely used hash function with 16-byte hash values, which is often used as a checksum of file downloads and suffices for creating collision-free encodings of instructions in our abstract dominator trees. The MD5 algorithm is the default hashing algorithm for fingerprints in the StoneDetector platform, and its application for encoding the description sets of the example in Fig. LABEL:fig2 is shown in the lower part of the same figure.
2.5 Comparison of Description Sets
Choosing one of the metrics in combination with a hashing algorithm, as presented in Sect. 2.3 and Sect. 2.4, for the distance function $\delta$ allows us to compare description sets and subsequently assess the similarity of code fragments. To this end, given two code fragments with respective description sets $D_{\mathit{set}}$ and $D^{\prime}_{\mathit{set}}$, the most similar path $q$ of $D^{\prime}_{\mathit{set}}$ is identified for each path $p$ of $D_{\mathit{set}}$, such that $q$ has a minimal value $\delta(p,q)$, and the arithmetic mean for these values over all paths $p\in D_{\mathit{set}}$ is computed. Formally, we can define the distance of the two description sets $D_{\mathit{set}},D^{\prime}_{\mathit{set}}\in\mathbb{D_{\mathit{set}}}$, having $|D_{\mathit{set}}|\leq|D^{\prime}_{\mathit{set}}|$ without loss of generality, as a function $\Delta$ such that:
3 Architecture and Implementation
@@
The approach to code clone detection introduced in Sect. 2 has been implemented in the StoneDetector platform [21, 33] thus providing a full-scale stand-alone clone detector for Java. As input, StoneDetector expects a number of Java source or Bytecode files for which the platform generates a list of all code clone pairs, which are identified and mapped back to the original code by file location and line numbers. The resulting information can then be used, for example, to search for similar programs in code repositories [34], to analyze the evolution of code [35], or to track bugs and vulnerabilities propagated via duplicated code [36].
Figure 5: Overview of the StoneDetector platform
@@
An overview of the StoneDetector platform can be seen in Fig. 5, which shows its inner structure and how its components interact with each other. Considering, to begin with, Java source code, the platform starts with performing lexical and syntax analysis on the given Java source files. This step yields the abstract syntax trees, which are subsequently subject to control flow analysis for generating (abstract) dominator trees. Afterward, paths are extracted from the generated dominator trees and linearized into description sets. In case fingerprints are enabled, the individual nodes of the description sets are further encoded using a hashing algorithm, which yields fingerprint sets. Either raw description sets or fingerprint sets are eventually transferred to the clone matcher, which checks for similarities using one out of the several available string metrics. Based upon this similarity analysis, code clones are identified and returned for further processing. We will next provide further details on the individual components of the described processing pipeline for Java source code files:
@@
The lexical and syntax analysis in the Java source code frontend of the StoneDetector platform is utilizing the Spoon analysis framework [37] [^2] https://spoon.gforge.inria.fr/. The Spoon framework includes libraries for parsing Java source code files up to Java version 20 into abstract syntax trees (AST) and provides the AST metamodel and functionality for manipulating them.
Control Flow/Dominator Analysis
The Spoon framework also supports generating control flow graphs of individual Java functions based on its AST model. We can therefore resort to Spoon’s control flow builder for our control flow analysis. However, the framework currently includes neither a meta-model for dominator trees nor any dominator analysis. For the construction of dominator trees, the respective builder from the WALA framework [38] [^3] https://github.com/wala/WALA has been integrated into the StoneDetector platform. As before, dominator trees are built at the function level. In addition, due to the coarse abstraction for Java’s exception handling in control flow graphs generated by Spoon, we have implemented a standard interprocedural exception analysis for Java as described by Chang et al. [39]. Using the analysis allows for a more fine-grained control flow mapping of Java’s exception handling. The analysis is, though, optional, but experiments have shown that enabling exception analysis implies better results for clone detection.
@@
As described in Sect. 2, dominator trees are transformed into description sets in order to subsequently perform textual clone matching. To this end, a (abstract) dominator tree is traversed such that the set of its paths starting at the root node and ending at leaf nodes, respectively, is generated (cf. Fig. LABEL:fig2). Note that we in general perform identifier and literal abstraction in the resulting description sets, much like as in other clone detectors, e.g., NiCad [9]. Though function names at call sites may optionally be conserved, which provided better experimental results in some cases. We therefore offer a switch for enabling/disabling identifier abstraction for function names at call sites. Likewise, the StoneDetector platform treats encoding of description sets into fingerprint sets via a hashing function as optional. Raw description sets can thus be directly committed to the clone matcher or beforehand transformed such that the contained instructions are encoded using one of the several implemented hashing algorithms, i.e., MD5, 4-byte prime number hashing, or LSH.
@@
The clone matching component receives a set of description or fingerprint sets as input and performs pairwise similarity checks. As described in Sect. 2, the similarity of description sets, i.e., code fragments, is traced back to the accumulated similarity of its paths in our approach. Path similarity is computed using one out of several string metrics available in the StoneDetector platform. Among the currently available metrics are the Hamming distance [25], both the normal and the weighted Levenshtein distance [26], the Needleman-Wunsch-based metric [27], as well as the original or modified LCS-based metric [28]. Further supported metrics include the N-Gram, Jaccard, cosine, Jaco-Winkler, and Damerau-Levenshtein metrics.
@@
The thus defined processing pipeline in the StoneDetector platform is parallelized. On the one hand, generation of description sets can be done in parallel with the help of streams. If all description sets have been eventually generated, clone matching begins. This process, on the other hand, can also be done in parallel, again using streams. The configuration of a specific run of the StoneDetector platform allows for the specification of the number of threads to be used in both computations. In order to further increase throughput, especially when processing larger numbers of code fragments, a copy strategy can optionally be used for description set handling. NiCad uses a similar technique for improving its runtime behavior [9, 14]. In StoneDetector, however, the technique has been adopted in such a way that the number of identified clones remains unchanged whether or not the technique is enabled. The basic idea of the copy strategy is to continue with further comparisons only for one of the two code fragments $f$ and $g$ whenever a clone pair $f$ and $g$ is found: Suppose comparisons are continued for fragment $f$ and terminated for fragment $g$. After all comparisons are finished, code clones found for $f$ are added to the clones already found for $g$, up to termination. Note that in contrast to NiCad’s implementation, in our variant of the copy strategy, two code fragments are only considered as a clone pair if they are identical to each other, i.e., if they have a distance value of zero. Using our copy strategy, runtime behavior does not improve as much as in NiCad, however, accuracy of the resulting code clones is guaranteed.

### v2

@@ Abstract
Copy & paste is a widespread practice when developing software and, thus, duplicated and subsequently modified code occurs frequently in software projects. Since such code clones, i.e., identical or similar fragments of code, can bloat software projects and cause issues like bug or vulnerability propagation, their identification is of importance. In this paper, we present StoneDetector and its underlying method for finding code clones in Java source and Bytecode. StoneDetector implements a conventional clone detection approach based upon the textual comparison of paths derived from the code’s representation by dominator trees. In this way, the tool does not only find exact and syntactically similar near-miss code clones, but also code clones that are harder to detect due to their larger variety in the syntax. We demonstrate StoneDetector’s versatility as a conventional clone detection tool and analyze its various available configuration parameters, including the usage of different string metrics, hashing algorithms, etc. In our exhaustive evaluation with other conventional clone detectors on several state-of-the-art benchmarks, we can show StoneDetector’s performance and scalability in finding code clones in both, Java source and Bytecode.
keywords:
code clone , clone detection , Java , static analysis
journal: Journal of Systems and Software
@@
[dhge]organization=Cooperative University Gera-Eisenach,addressline=Weg der Freundschaft 4, city=Gera, postcode=07546, country=Germany
@@
[fsu]organization=Friedrich Schiller University Jena,addressline=Ernst-Abbe-Platz 2, city=Jena, postcode=07743, country=Germany
@@ 1 Introduction
Code clones [83] are identical or similar fragments of code that, e.g., arise due to copy-and-paste practices in software development. Since such practices are quite common, code clones can be frequently found, as empirical studies on code duplication have shown in the past [84, 123, 64]. While code cloning may initially speed up software development or allow for reusing working code, it also introduces long-term challenges: Duplicated code inflates code bases and increases the maintenance burden, as developers have to manually track and update all copies for fixing bugs or implementing code changes. The effects of code cloning are thus controversial [45, 43] and detecting code clones is a vital step toward maintaining code quality, reducing technical debt, and improving program comprehension.
@@ 10 FileInputStream fis = new FileInputStream(f);
The presented clone detection method is implemented in StoneDetector, which is publicly available [^1] https://stonedetector.fmi.uni-jena.de/ and allows for finding code clones in Java source and Bytecode. Using StoneDetector enabled us to perform an in-depth evaluation of the method’s performance on various state-of-the-art benchmarks, including BigCloneBench [102, 103], Google Code Jam, Project CodeNet [78], and GPTCloneBench [4]. As a result, we found that our clone detection method and its implementation in StoneDetector allow for finding exact and near-miss code clones in Java source code at competitive performance with other conventional code clone detectors, including NiCad [22], iClones [31], SourcererCC [88], Deckard [39], CCAligner [110], NIL [69], CloneWorks [104] and Oreo [87]. Additionally, StoneDetector in particular excels at finding source code clones with larger syntactical variance at high precision. StoneDetector also scales to large codebases comprising multiple 100 million lines of code and features runtimes comparable to the state of the art. Furthermore, we have experimented with different configuration parameters for clone detection, e.g., threshold value, code clone size, applied string metrics, and hashing algorithms, etc., and thereby show the versatility of StoneDetector besides identifying its optimal configuration. Eventually, we analyzed the method’s potential in also finding code clones in compilation artifacts like Java Bytecode, in cases where the source code is unavailable in the first place. In summary, the contributions of this paper are:
@@
    We demonstrate the feasibility of finding code clones in Java source code fragments by encoding and textually comparing paths of the code fragments’ dominator trees.
@@
    We provide the publicly available tool StoneDetector with an implementation of our code clone detection method as a versatile conventional code clone detector and evaluate its performance and scalability.
@@
This work is based upon the authors’ prior conference paper on code clone detection presented at the 2021 IEEE International Conference on Software Maintenance and Evolution [8] and extends it in various ways. First, an analysis of the usage of different string metrics within StoneDetector has been added to the paper, including distance metrics like Levenshtein, Needleman-Wunsch, Hamming distance, and longest common subsequence (LCS)-based metric. Among others, the analysis revealed the versatility of StoneDetector as a clone detection framework: For example, using LCS provided the best tradeoff in terms of recall, precision, and scalability. In contrast, using the Hamming distance yielded a very fast clone detector with impaired performance in finding harder-to-detect Type 3/4 code clones. Second, we now derive the optimal configuration of StoneDetector for clone detection by means of an in-depth empirical analysis of its several configuration parameters (threshold value for path comparison, etc.). Furthermore, expanding the approach for detecting large-gap and subclones, as presented in our previous conference paper [8], we added locality-sensitive hashing (LSH) and a modified LCS metric to StoneDetector, which resulted in performance gains for these types of code clones. We now also include a more exhaustive analysis of StoneDetector’s performance in finding harder-to-detect Type 3/4 code clones by including three more state-of-the-art benchmarks for code clone detection besides BigCloneBench, i.e., Google Code Jam, Project CodeNet, and GPTCloneBench, and three more clone detectors, i.e., iClones, CloneWorks [104], and NIL, in the evaluation. Moreover, we also extended our previous approach in Amme et al. [8] for clone detection on stack-based and register-based Java Bytecode representations, such that StoneDetector can be used for both finding code clones in Java source code and Bytecode.
@@
The rest of the paper is structured as follows: Sect. 2 introduces StoneDetector’s overall approach and architecture. The underlying clone detection method based on encodings of paths in the dominator tree and string metrics for path comparison are described in detail in Sect. 3. In Sect. 4, we present the in-depth evaluation of our approach on various benchmarks for Java code clone detection, including the comparison with state-of-the-art clone detectors and the thorough analysis of StoneDetector’s various configuration parameters. Related work on code clone detection is discussed in Sect. 5 and Sect. 6 finally concludes the paper and mentions future work.
2 Overall Approach and Architecture
@@
Our approach to code clone detection has been implemented in StoneDetector [8, 93] thus providing a full-scale stand-alone clone detector for Java. As input, StoneDetector expects a number of Java source or Bytecode files for which the tool generates a list of all code clone pairs, which are identified and mapped back to the original code by file location and line numbers. The resulting information can then be used, for example, to search for similar programs in code repositories [47], to analyze the evolution of code [72], or to track bugs and vulnerabilities propagated via duplicated code [61].
Figure 3: Overview of the tool StoneDetector
@@
An overview of StoneDetector can be seen in Fig. 3, which shows its inner structure and how its components interact with each other. Considering, to begin with, Java source code, the tool starts with performing lexical and syntax analysis on the given Java source files. This step yields the abstract syntax trees, which are subsequently subject to control flow analysis for generating (abstract) dominator trees. Afterward, paths are extracted from the generated dominator trees and linearized into description sets. The individual nodes of the description sets can be further encoded using a hashing algorithm, which yields fingerprint sets. Either raw description sets or fingerprint sets are eventually transferred to the clone matcher, which checks for similarities using one out of the several available string metrics. Based upon this similarity analysis, code clones are identified and returned for further processing. We will next provide more information on the individual components of the described processing pipeline for Java source code files. For a more elaborative introduction of the clone detection method based on dominator tree and string metrics we refer the reader to Sect. 3.
@@
The lexical and syntax analysis in the Java source code frontend of StoneDetector is utilizing the Spoon analysis framework [76] [^2] https://spoon.gforge.inria.fr/. The Spoon framework includes libraries for parsing Java source code files up to Java version 20 into abstract syntax trees (AST) and provides the AST metamodel and functionality for manipulating them.
Control/Dominator Analysis
The Spoon framework also supports generating control flow graphs of individual Java functions based on its AST model. We can therefore resort to Spoon’s control flow builder for our control flow analysis. However, the framework currently includes neither a meta-model for dominator trees nor any dominator analysis. For the construction of dominator trees, the respective builder from the WALA framework [24] [^3] https://github.com/wala/WALA has been integrated into StoneDetector. As before, dominator trees are built at the function level. In addition, due to the coarse abstraction for Java’s exception handling in control flow graphs generated by Spoon, we have implemented a standard interprocedural exception analysis for Java as described by Chang et al. [17]. Using the analysis allows for a more fine-grained control flow mapping of Java’s exception handlin and experiments have shown that enabling exception analysis implies better results for clone detection.
@@
Dominator trees are transformed into description sets in order to subsequently perform textual clone matching. To this end, a (abstract) dominator tree is traversed such that the set of its paths starting at the root node and ending at a leaf nodes, respectively, is generated. Note that we in general perform identifier and literal abstraction in the resulting description sets, much like as in other clone detectors, e.g., NiCad [85]. Though function names at call sites may optionally be conserved, which provided better experimental results in some cases. We therefore offer a switch for enabling/disabling identifier abstraction for function names at call sites. Likewise, StoneDetector treats encoding of description sets into fingerprint sets via a hashing function as optional. Raw description sets can thus be directly committed to the clone matcher or beforehand transformed such that the contained instructions are encoded using one of the several implemented hashing algorithms, i.e., MD5, 4-byte prime number hashing, or Locality-Sensitive Hashing (LSH).
@@
The clone matching component receives a set of description or fingerprint sets as input and performs pairwise similarity checks. The similarity of description sets, i.e., code fragments, is traced back to the accumulated similarity of its paths in our approach. Path similarity is computed using one out of several string metrics available in StoneDetector. Among the currently available metrics are the Hamming distance [34], the Levenshtein distance [56], the Needleman-Wunsch-based metric [71], as well as the LCS-based metric [12]. Further supported metrics include the N-Gram, Jaccard, Cosine, Jaco-Winkler, and Damerau-Levenshtein metrics.
@@
The thus defined processing pipeline in StoneDetector is parallelized. On the one hand, generation of description sets can be done in parallel with the help of streams. If all description sets have been eventually generated, clone matching begins. This process, on the other hand, can also be done in parallel, again using streams. The configuration of a specific run of StoneDetector allows for the specification of the number of threads to be used in both computations. In order to further increase throughput, especially when processing larger numbers of code fragments, a copy strategy can optionally be used for description set handling. NiCad uses a similar technique for improving its runtime behavior [85, 22]. In StoneDetector, however, the technique has been adopted in such a way that the number of identified clones remains unchanged whether or not the technique is enabled. The basic idea of the copy strategy is to continue with further comparisons only for one of the two code fragments $f$ and $g$ whenever a clone pair $f$ and $g$ is found: Suppose comparisons are continued for fragment $f$ and terminated for fragment $g$. After all comparisons are finished, code clones found for $f$ are added to the clones already found for $g$, up to termination. Note that in our variant of the copy strategy, two code fragments are only considered as a clone pair if they are identical to each other, i.e., if they have a distance value of zero. Using our copy strategy, runtime behavior does not improve as much as in NiCad, however, accuracy of the resulting code clones is guaranteed.